I am trying to solve this equation: 19 - 3n = negative 2n. I do not understand the steps needed to solve this problem.
Thank You,

Answer by checkley71(8403)

19-3N=-2N OR 19=3N-2N OR 19=N
PROOF 19-3*19=-2*19 OR 19-57=-38 OR -38=-38